# Addressing Model and Data Heterogeneity in Multimodal Large Language Model Training

### v1

@@ Abstract
We present MMScale, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters. MMScale exploits the system characteristics of multimodal LLM training to achieve high efficiency and scalability. The core of MMScale is the adaptive resource allocation and data-aware reordering techniques to eliminate the model and data heterogeneity respectively. We also tailor system optimizations for multimodal LLM training to offload certain operations from the GPU training. We evaluate MMScale across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs. The experimental results show that MMScale achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput. The ablation study shows the main techniques of MMScale are both effective and lightweight.
@@ 1 Introduction
Megatron-LM [17] is the state-of-the-art framework for training LLMs, which comprise a stack of homogeneous transformer layers. It can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM. However, training multimodal LLMs with Megatron-LM poses two substantial challenges: model heterogeneity and data heterogeneity. The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity. The model heterogeneity across different modules (i.e., modality encoder, LLM backbone, and modality generator) introduces severe pipeline bubbles, resulting in poor GPU utilization. Meanwhile, data heterogeneity bursts onto the scene due to the intricate and unstructured nature of multimodal input data. The data heterogeneity across different modality data leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles. These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§7.1).
@@
To this end, we present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs. MMScale achieves state-of-the-art MFU which is close to unimodal (i.e., text-only) LLM training and effectively scales to large clusters with thousands of GPUs. The core of MMScale is adaptive resource allocation and data-aware reordering. These two techniques leverage the distinctive system characteristics of multimodal LLM training to eliminate model heterogeneity and data heterogeneity respectively.
@@
For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2). Based on this analysis, we propose disaggregated training, which tailors the resource allocation and parallelism strategies across the modality encoder, LLM backbone, and modality generator. We also develop a novel and lightweight algorithm to navigate this complicated design space to choose the optimal resource and parallelism configurations. This innovative approach minimizes the pipeline bubbles and achieves optimal training efficiency.
@@
For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3) and introduce tailored algorithms that strategically reorder training data while maintaining synchronous training semantics. The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups. The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles. This two-level approach effectively mitigates training stragglers caused by data heterogeneity.
@@
In addition, we customize system optimizations for multimodal LLM training to offload certain operations from the critical path (i.e., GPU computation). We implement an asynchronous data loader to pre-process the training data with negligible runtime overhead. We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
@@
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity. It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
@@
We implement disaggregated training across the three modules in multimodal LLMs. We propose adaptive resource allocation and data-aware reordering to effectively address model and data heterogeneity, respectively.
We implement MMScale and conduct experiments on our production cluster with thousands of GPUs. The experimental results show that MMScale achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
@@ 2.1 Multimodal Large Language Model Training
Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator [24, 15, 16]. The modality encoder transforms input data from different modalities (e.g., ViT [25] for images and Beats [26] for audios) into an intermediate representation (i.e., an embedding tensor), which is then projected into a unified embedding space across modalities with input projection layers (e.g., MLP and cross-attention). The LLM backbone, typically a transformer model (e.g., GPT [19, 20] and Llama [22]), processes the multimodal embeddings to discern the intricate data patterns and inter-modal relationships. The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality. Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs to generate coherent and contextually relevant results.
@@
However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid resource allocation method (i.e., the multimodal modules share the same DP and TP strategies with LLM backbone). It only achieves ~20% MFU (§7.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
@@ 3 MMScale Overview
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity. MMScale exploits the characteristics of parallelism training of multimodal LLMs to achieve high efficiency and scalability. MMScale eliminates the model heterogeneity by adaptive resource allocation (§4) and harnesses the data heterogeneity by data-aware reordering (§5). In addition, MMScale adopts some system optimizations tailored for multimodal LLM training (§6). Here we provide a brief overview of MMScale as Figure 8 shows.
@@
MMScale orchestrator. Before training, MMScale employs a training orchestrator to determine the resource allocation and parallelism strategy for each module in multimodal LLMs. This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution. Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time. Based on the profiling results, the training orchestrator decides the optimal resource allocation and parallelism strategy with adaptive resource allocation for one specific training task, as detailed in §4.
@@
MMScale initializer. Then, MMScale initializes the parallelism units for the modality encoder, LLM backbone, and modality generator, respectively. MMScale allocates different numbers of GPUs to each unit. Each unit then establishes the communication group across different nodes and GPUs. The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states. Finally, MMScale conducts several communication trials to warm up the system and test connectivity.
@@
MMScale runtime. After initialization, MMScale begins the main training process. In each training iteration, it first retrieves a global batch of training samples from the distributed file system for pre-processing. Then, MMScale performs data-aware reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31]. This module effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §5. The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline. Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33]. Additionally, MMScale adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
@@ 4 Addressing Model Heterogeneity
Before diving into the details, we first introduce the disaggregation of the training across the three modules (i.e., modality encoder, LLM backbone, and modality generator) in multimodal LLMs. This disaggregation facilitates adaptive GPU resource allocation among the different modules based on the training workload. Subsequently, we formulate the problem of adaptive resource allocation to minimize the training time per iteration. Finally, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
4.1 Disaggregated Training
Disaggregating the training among the three modules is the foundation of adaptive resource allocation in MMScale. Figure 9 demonstrates the training topology of disaggregated training in MMScale. Different from the rigid resource allocation in Megatron-LM (i.e., Figure 2), MMScale is able to adaptively adjust the resource allocation. For instance, MMScale allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator. Additionally, the projector layers are colocated with either the modality encoder or generator, with their number of replicas adapting as needed. We implement the disaggregated training through a dedicated module, i.e., parallelism unit.
Parallelism unit. At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy. MMScale introduces a module, parallelism unit, composed of one or more PP stages. Each unit can adopt its own DP and TP strategies and form a specific communication group. Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units. Users are only required to specify the DP and TP configurations for each parallelism unit, and MMScale automatically sets up the communication group and communication broker. MMScale treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units. The detailed implementation of parallelism unit is discussed in §6.
@@ 4.2 Problem Formulation
With disaggregated training, we are able to adaptively allocate GPU resources across the three modules. The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration. Exhaust search is infeasible due to the large search space of the parallelism strategy and cluster size. One strawman solution is to allocate the GPU resources proportional to the model flops of each module. However, this method falls short as it overlooks complex patterns in parallelism training. For instance, different DP and TP sizes in modality encoder could lead to different computation times while the model flops are the same. Before diving into the adaptive resource allocation algorithm, we first formulate the optimization problem.
@@
Modality encoder and generator. In MMScale, the modality encoder is regarded as a parallelism unit with PP size . Let the TP size be and the DP size be . The number of GPUs allocated to modality encoder is . The microbatch size is which is determined by the LLM backbone. Let the forward time (including communication time) of the entire modality encoder be . The forward time of one PP stage for one microbatch in the modality encoder is . Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality encoder.
@@
Therefore, the objective function is to minimize . For the backward pass, the objective function remains analogous to that of the forward pass. Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time. This formulation holds for GPipe and 1F1B. We will retrofit the formulation to adapt to VPP later. As for the communication time, it consists of the time of P2P communication in PP and collective communication in DP. It is modeled as the communication volume divided by the communication bandwidth. If the sizes of TP and DP are fixed, the communication time remains constant.
Constraints. Besides the objective function, we must consider constraints to ensure training feasibility. The first constraint is the resource constraint. The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster. The second constraint involves memory. We consider LLM backbone. Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states. The memory of model parameters and gradients on one GPU is calculated as: , where denotes total memory for the LLM parameters and gradients. The memory for optimizer states on one GPU (with ZeRO-1 optimization) is: , where denotes the total memory for the optimizer states. ZeRO-1 partitions the optimizer states across DP groups. The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch of activation states across the entire LLM. In 1F1B, the first PP stage requires storage for microbatches of activation states. We eschew using GPipe in MMScale since GPipe consumes more memory. The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity. As for modality encoder and generator, the formulation is similar.
4.3 Adaptive Resource Allocation
@@
Convex optimization on resource allocation. Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., the resource allocation). We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups. The PP size of the LLM backbone is calculated as . The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of the TP and DP size. This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems. In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive. Therefore, the objective function is convex. Similarly, the constraint functions are also convex. As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35]. The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy to minimize iteration time.

### v2

# DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models

@@ Abstract
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters. The core of DistTrain is the disaggregated training technique that exploits the characteristics of multimodal LLM training to achieve high efficiency and scalability. Specifically, it leverages disaggregated model orchestration and disaggregated data reordering to address model and data heterogeneity respectively. We also tailor system optimization for multimodal LLM training to overlap GPU communication and computation. We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs. The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput. The ablation study shows the main techniques of DistTrain are both effective and lightweight.
@@ 1 Introduction
Existing LLM training framework, e.g., Megatron-LM [17], can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM. However, training multimodal LLMs poses two substantial challenges: model heterogeneity and data heterogeneity. The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity. The model heterogeneity across different modules (i.e., modality encoder, LLM backbone, and modality generator) introduces severe pipeline bubbles, resulting in poor GPU utilization. Meanwhile, data heterogeneity bursts onto the scene due to the intricate and unstructured nature of multimodal input data. The data heterogeneity across different modality data leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles. These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§8.1).
@@
To this end, we present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs. DistTrain achieves state-of-the-art MFU which is close to unimodal LLM training and effectively scales to large clusters with thousands of GPUs. The core principle of DistTrain is disaggregated training, including GPU training disaggregation and CPU preprocessing disaggregation, which facilitate to address model and data heterogeneity respectively.
@@
For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2). The GPU training disaggregation of modality encoder, LLM backbone, and modality generator enables adaptive orchestration across the three modules in multimodal LLM. Building on this, we propose disaggregated model orchestration that navigates the complicated design space to choose the optimal resource and parallelism configurations. This innovative approach minimizes the pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
@@
For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3). The CPU preprocessing disaggregation allows efficient data preprocessing (e.g., data decompression and reordering) with negligible runtime overhead. From this foundation, we incorporate disaggregated data reordering into the preprocessing to strategically reorder training data without additional overhead. The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups. The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles. This two-level approach effectively mitigates data heterogeneity.
@@
In addition, we customize system optimization for multimodal LLM training. We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
@@
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity. It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
@@
We propose disaggregated training for multimodal LLM. It leverages disaggregated model orchestration and disaggregated data reordering to effectively address model and data heterogeneity, respectively.
We implement DistTrain and conduct experiments on our production cluster with thousands of GPUs. The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
@@ 2.1 Multimodal Large Language Model Training
Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator [24, 15, 16]. The modality encoder transforms input data from different modalities (e.g., ViT [25] for images and Beats [26] for audios) into an intermediate representation (i.e., an embedding tensor), which is then projected into a unified embedding space across modalities with input projection layers (e.g., MLP and cross-attention). The LLM backbone, typically a transformer model (e.g., GPT [19, 20] and Llama [22]), processes the multimodal embeddings to discern the intricate data patterns and inter-modal relationships. The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality. Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs for generation.
@@
However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid model orchestration method. (i.e., the multimodal modules share the same DP and TP strategies with LLM backbone). It only achieves ~20% MFU (§8.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
@@ 3 DistTrain Overview
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity. DistTrain proposes disaggregated training for multimodal LLM to achieve high efficiency and scalability. DistTrain eliminates the model heterogeneity by disaggregated model orchestration (§5) and harnesses the data heterogeneity by disaggregated data reordering (§6). In addition, DistTrain adopts some system optimizations tailored for multimodal LLM training (§7). Here we provide a brief overview of DistTrain as Figure 8 shows.
@@
DistTrain manager. Before training, DistTrain employs a training manager to determine the resource allocation and parallelism strategy for each module in multimodal LLMs. This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution. Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time. Based on the profiling results, the training manager decides the optimal resource allocation and parallelism strategy with disaggregated model orchestration for one specific training task, as detailed in §5.
@@
DistTrain initializer. DistTrain then initializes the GPU training disaggregation for modality encoder, LLM backbone, and modality generator, respectively. DistTrain allocates different numbers of GPUs to each parallelism unit. Each unit then establishes the specific communication group. The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states. Finally, DistTrain conducts several communication trials to warm up the system and test connectivity.
@@
DistTrain runtime. At runtime, the dedicated CPU nodes (i.e., CPU preprocessing disaggregation) retrieve training samples from the distributed file system for preprocessing. It performs disaggregated data reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31]. This reordering effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §6. In each iteration, the main training process receives the preprocessed data asynchronously from the CPU nodes. The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline. Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33]. Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
4 Disaggregated Training
To address the model and data heterogeneity in multimodal LLM training, we first introduce DistTrain’s core principle: disaggregated training. It includes GPU training disaggregation and CPU preprocessing disaggregation. GPU training disaggregation provides opportunities for adaptive model orchestration across the three modules to address the model heterogeneity. CPU preprocessing disaggregation facilitates data preprocessing to address data heterogeneity with negligible runtime overhead.
4.1 GPU Training Disaggregation
Figure 9 demonstrates the training topology of disaggregated GPU training in DistTrain. Different from the rigid model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy. For instance, DistTrain allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator. Additionally, the projector layers are co-located with either the modality encoder or generator, with their number of replicas adapting as needed. We implement GPU training disaggregation through a dedicated module, i.e., parallelism unit.
Parallelism unit. At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy. DistTrain introduces a module, parallelism unit, composed of one or more PP stages. Each unit can adopt its own DP and TP strategies and form a specific communication group. Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units. Users are only required to specify the DP and TP configurations for each parallelism unit, and DistTrain automatically sets up the communication group and communication broker. DistTrain treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units. The detailed implementation of parallelism unit is discussed in §7.
4.2 CPU Preprocessing Disaggregation
When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data. The latter significantly increases data preprocessing time. For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images. The text is just kilobytes, whereas the images are total of 120 megabytes. Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds and interfere with the co-located training process. DistTrain disaggregates the CPU data preprocessing from the GPU training process with a producer-consumer model. The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously with the GPU training process. The consumer, i.e., the main training process, receives this preprocessed data for training. The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available. This disaggregation guarantees that the CPU preprocessing does not interfere with the GPU training process and achieves negligible data preprocessing overhead.
@@ 5 Addressing Model Heterogeneity
Disaggregated training enables disaggregated model orchestration among the different modules based on the training workload. We first formulate the problem of disaggregated model orchestration to minimize the training time per iteration. Then, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
@@ 5.1 Problem Formulation
With disaggregated training, we are able to adaptively orchestrate the three modules. The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration. Exhaust search is infeasible due to the large search space, particularly in large cluster. One strawman solution is to allocate the resources proportional to the model flops of each module. However, this method falls short as it overlooks complex patterns in parallelism training. Before diving into disaggregated model orchestration, we first formulate the optimization problem.
@@
Modality encoder and generator. In DistTrain, the modality encoder is regarded as a parallelism unit with PP size . Let the TP size be and the DP size be . The number of GPUs allocated to modality encoder is . The microbatch size is which is determined by the LLM backbone. Let the forward time (including communication time) of the entire modality encoder be . The forward time of one PP stage for one microbatch in the modality encoder is . Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality generator.
@@
Therefore, the objective function is to minimize . For the backward pass, the objective function remains analogous to that of the forward pass. Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time. This formulation holds for GPipe and 1F1B. We will retrofit the formulation to adapt to VPP later. TP communication is incorporated into the functions , and , which are calibrated through interpolation from actual trials. The communication time of DP and PP is modeled as the communication volume divided by the bandwidth. If the DP sizes are fixed, the DP communication time remains constant. For PP, the communication time equals the PP size multiplied by the communication time of a single layer’s output tensor.
Constraints. Besides the objective function, we must consider constraints to ensure training feasibility. The first constraint is the resource constraint. The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster. The second constraint involves memory. We consider LLM backbone. Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states. The memory of model parameters and gradients on one GPU is calculated as: , where denotes total memory for the LLM parameters and gradients. The memory for optimizer states on one GPU (with ZeRO-1 optimization) is: , where denotes the total memory for the optimizer states. ZeRO-1 partitions the optimizer states across DP groups. The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch of activation states across the entire LLM. In 1F1B, the first PP stage requires storage for microbatches of activation states. We eschew using GPipe in DistTrain since GPipe consumes more memory. The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity. As for modality encoder and generator, the formulation is similar.
5.2 Disaggregated Model Orchestration
@@
Convex optimization. Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., resource allocation). We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups. The PP size of the LLM backbone is calculated as . The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of TP and DP size. This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems. In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive. Therefore, the objective function is convex. Similarly, the constraint functions are also convex. As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35]. The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy.